# Entropy-Wasserstein regularization, defective local concentration and a cutoff criterion beyond non-negative curvature

### v1

@@ Abstract
Notions of positive curvature have been shown to imply many remarkable properties for Markov processes, in terms, e.g., of regularization effects, functional inequalities, mixing time bounds and, more recently, the cutoff phenomenon. In this work, we are interested in a relaxed variant of Ollivier’s coarse Ricci curvature, where a Markov kernel $P$ satisfies only a weaker Wasserstein bound $W_{p}(\mu P,\nu P)\leq KW_{p}(\mu,\nu)+M$ for constants $M\geq 0,K\in[0,1],p\geq 1$. Under appropriate additional assumptions on the one-step transition measures $\delta_{x}P$, we establish (i) a form of local concentration, given by a defective Talagrand inequality, and (ii) an entropy-transport regularization effect. We consider as illustrative examples the Langevin dynamics and the Proximal Sampler when the target measure is a log-Lipschitz perturbation of a log-concave measure. As an application of the above results, we derive criteria for the occurrence of the cutoff phenomenon in some negatively curved settings.
@@ 1.4 Negatively curved Markov chains
Having non-negative curvature is however a restrictive condition for a Markov chain, and the main motivation of this work is to investigate whether it is possible to relax Ollivier’s condition (1.1) while still deriving analogous consequences. In particular, our focus in this work is on Markov chains satisfying the following assumption.
@@ 1.5 Main results and paper organization
The first question that we study in Section 2 are the concentration of measure properties of the iterates $\mu_{k}=\operatorname{law}(X_{k})$ of the Markov chain at a finite-time, in the spirit of the local Poincaré inequality (1.6), but under the weaker Assumption 1.1. It turns out that, under the corresponding assumption on the initialization $\mu_{0}$ and the one-step transition measures $\delta_{x}P$, the distribution $\mu_{k}$ satisfies a defective version of Talagrand’s transport-entropy inequality, cf. Theorem 2.3. For example, for the Langevin dynamics under Assumption 1.3, we deduce that for all $\mu\in\mathcal{P}\left(\mathbb{R}^{d}\right),x\in\mathbb{R}^{d},t>0$
@@ 2 Defective local concentration
Our focus in this work is more generally on local functional inequalities and the relationship with Assumption 1.1, and in this sense our main result in this section is Theorem 2.3 below. Inspired by earlier work of Marton [Mar96] and the subsequent developments in [DGW04, ELL17], which were however concerned with the classical non-defective transport inequalities, we make the key assumption that the one-step transition kernel satisfies a defective transport-entropy inequality:
@@
Under Assumptions 1.1 and 2.2 (for the same $p$), if $\mu_{0}\in\operatorname{{T_{p}}H}\left(J,S\right)$ then

### v2

@@ Abstract
Notions of positive curvature have been shown to imply many remarkable properties for Markov processes, in terms, e.g., of regularization effects, functional inequalities, mixing time bounds, and, more recently, the cutoff phenomenon. In this work, we are interested in a relaxed variant of Ollivier’s coarse Ricci curvature, where a Markov kernel $P$ satisfies only a weaker Wasserstein bound $W_{p}(\mu P,\nu P)\leq KW_{p}(\mu,\nu)+M$ for constants $M\geq 0,K\in[0,1],p\geq 1$. Under appropriate additional assumptions on the one-step transition measures $\delta_{x}P$, we establish (i) a form of local concentration, given by a defective Talagrand inequality, and (ii) an entropy-transport regularization effect. We consider as illustrative examples the Langevin dynamics and the Proximal Sampler when the target measure is a log-Lipschitz perturbation of a log-concave measure. As an application of the above results, we derive criteria for the occurrence of the cutoff phenomenon in some negatively curved settings.
@@ 1.4 Negatively curved Markov chains
Having non-negative curvature is, however, a restrictive condition for a Markov chain, and the main motivation of this work is to investigate whether it is possible to relax Ollivier’s condition (1.1) while still deriving analogous consequences. In particular, our focus in this work is on Markov chains satisfying Assumption 1.1 below: this assumption was also briefly discussed in Ollivier’s paper [Oll09], where it was referred to as “positive curvature up to MM”. While the motivation in [Oll09] was topological, i.e., to turn positive curvature into an open property, in his subsequent survey Ollivier explicitly raised the question of understanding which consequences of (1.1) extend to the more general setting of (1.11) below, see [Oll10, Problem O].
@@ 1.5 Main results and paper organization
The first question that we study in Section 2 is concerned with the concentration of measure properties of the iterates $\mu_{k}=\operatorname{law}(X_{k})$ of the Markov chain at a finite-time, in the spirit of the local Poincaré inequality (1.6), but under the weaker Assumption 1.1. It turns out that, under the corresponding assumption on the initialization $\mu_{0}$ and the one-step transition measures $\delta_{x}P$, the distribution $\mu_{k}$ satisfies a defective version of Talagrand’s transport-entropy inequality, cf. Theorem 2.4. For example, for the Langevin dynamics under Assumption 1.3, we deduce that for all $\mu\in\mathcal{P}\left(\mathbb{R}^{d}\right),x\in\mathbb{R}^{d},t>0$
@@ 2 Defective local concentration
Remark 2.2.
This lemma already highlights some important differences between the defective form of the inequality and the standard one, when $p>1$. To see this, consider for $\varepsilon>0$ the probability measures $\pi_{\varepsilon},\mu\in\mathcal{P}(\mathbb{R})$ given by $\pi_{\varepsilon}=\frac{1}{2}(\delta_{-\varepsilon}+\delta_{\varepsilon})$ and $\mu=\delta_{0}$. Then, $W_{\infty}\left(\mu,\pi\right)\leq\varepsilon$ and $\mu\in\operatorname{{T_{p}}H}\left(0\right)$ for every $p\geq 1$, and so by Lemma 2.1 we deduce that $\pi\in\operatorname{{T_{p}}H}\left(0,2\varepsilon\right)$. However, for any $C>0$ and $p>1$ we necessarily have that $\pi\notin\operatorname{{T_{p}}H}\left(2C\right)$, since the support of $\pi$ is disconnected (cf. [FS18, Lemma 2.1], whose argument is easily adapted to the general case $p>1$); similarly, $\pi$ does also not satisfy other classical functional inequalities, such as the Poincaré inequality. In many situations, including in the study of the cutoff phenomenon in Section 4, it is still desirable to think of the measure $\pi_{\varepsilon}$ as being “well-concentrated” when $\varepsilon$ is small, and one important motivation for the introduction of the defective transport-entropy inequalities is to rigorously capture this intuition, despite the failure of some classical functional inequalities.
Our focus in this work is more generally on local functional inequalities and the relationship with Assumption 1.1, and in this sense, our main result in this section is Theorem 2.4 below. Inspired by earlier work of Marton [Mar96] and the subsequent developments in [DGW04, ELL17], which were, however, concerned with the classical non-defective transport inequalities, we make the key assumption that the one-step transition kernel satisfies a defective transport-entropy inequality:
@@
Under Assumptions 1.1 and 2.3 (for the same $p$), if $\mu_{0}\in\operatorname{{T_{p}}H}\left(J,S\right)$ then
@@ 4.2 Proximal sampler
Acknowledgments
The author thanks Sam Power, Justin Salez, and Peter Whalley for useful comments and references.